THE PERTURBATION OF C_{3} BY THE RARE GAS ATOM STUDIED BY THE SPECTROSCOPY OF THE C_{3} − AR/KR COMPLEX

GUIQIU ZHANG, INSTITUTE OF ATOMIC AND MOLECULAR SCIENCES, ACADEMIA SINICA, P. O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C. AND DEPARTMENT OF CHEMISTRY, XAN DONG TEACHER’S UNIVERSITY, PEOPLE REPUBLIC OF CHINA; \underline{YEN-CHU HSU}, INSTITUTE OF ATOMIC AND MOLECULAR SCIENCES, ACADEMIA SINICA, P. O. BOX 23-166, TAIPEI 107, TAIWAN, R. O. C. AND DEPARTMENT OF CHEM-ISTRY, NATIONAL TAIWAN UNIVERSITY, TAIPEI, TAIWAN, R. O. C..

The ground electronic state of the C_{3}Kr van der Waals complex has been studied by the wavelength-resolved emission from the Ã state. The vibrational levels of the C_{3}Kr complex have been compared with those of the C_{3}Ar to ensure the spectral assignments of these two complexes.The perturbation of the C_{3} bending vibration (both in-plane and out-of-plane bending motion) by the rare gas atom has been observed; the vibrational amplitude of the C_{3} bending motion of the complex becomes more restricted than the monomer.